OpenStudy (anonymous):

Angle C is an inscribed angle of circle P. Angle C measures (-3x - 6)° and arc AB measures (-4x)° . Find x.

OpenStudy (anonymous):



OpenStudy (nikato):

The inscribed angle is half the measure of its intercepted arc. So <C= 1/2 ( m. Arc AB) -3x-6=1/2 (-4x) Can you try to solve for x now?

OpenStudy (anonymous):

-6?

OpenStudy (raden):

yes, that's correct. x = -6

OpenStudy (anonymous):

thank you!